How many kilograms of coffee beans selling for $2.40 per kilogram should be mixed with 4 kilograms of coffee beans selling for $3.60 per kilogram to obtain a mixture selling for $2.88 per kilogram?
Answer by stanbon(75887)

How many kilograms of coffee beans selling for $2.40 per kilogram should be mixed with 4 kilograms of coffee beans selling for $3.60 per kilogram to obtain a mixture selling for $2.88 per kilogram?
------------
Amt. of $2.40 coffee = "x" kg ; Value of this is 2.4x dollars
Amt. of $3.60 coffee = 4 kg; Value of this is 3.6*4 = 14.4 dollars
Amt. of mixture = "x+4" kg; Value of this is 2.88(x+4) = (2.88x+ 11.52) dollars
-----------------
value + value = value
2.4x + 14.4 = 2.88x+11.52
-0.48x = -2.88
x= 6 kg (amt of $2.40 coffee to be added)
==========
Cheers,
Stan H.